### v2

# Learn2Synth: Learning Optimal Data Synthesis using Hypergradients for Brain Image Segmentation

Abstract
Domain randomization through synthesis is a powerful strategy to train networks that are unbiased with respect to the domain of the input images. Randomization allows networks to see a virtually infinite range of intensities and artifacts during training, thereby minimizing overfitting to appearance and maximizing generalization to unseen data. Although powerful, this approach relies on the accurate tuning of a large set of hyperparameters that govern the probabilistic distribution of the synthesized images. Instead of manually tuning these parameters, we introduce Learn2Synth, a novel procedure in which synthesis parameters are learned using a small set of real labeled data. Unlike methods that impose constraints to align synthetic data with real data (e.g., contrastive or adversarial techniques), which risk misaligning the image and its label map, we tune an augmentation engine such that a segmentation network trained on synthetic data has optimal accuracy when applied to real data. This approach allows the training procedure to benefit from real labeled examples, without ever using these real examples to train the segmentation network, which avoids biasing the network towards the properties of the training set. Specifically, we develop parametric and nonparametric strategies to enhance synthetic images in a way that improves the performance of the segmentation network. We demonstrate the effectiveness of this learning strategy on synthetic and real-world brain scans. Code is available at: https://github.com/HuXiaoling/Learn2Synth.
footnotetext: $\dagger$ Email: Xiaoling Hu (xihu3@mgh.harvard.edu); $\ddagger$ Co-senior authors
Figure 1: SynthSeg [6] (top left) is a domain randomization strategy that has shown great success in neuroimaging applications, where an image with random contrast is synthesized from a label map using a set of stochastic rules. This strategy contrasts with typical supervised training [56] (bottom left), where a model is trained on real labeled images (possibly augmented). Our proposed strategy, Learn2Synth (right), interleaves a typical SynthSeg step (1), that optimizes a segmentation network ($S_{\boldsymbol{\phi}}$) on synthetic data, with a supervised step (2). Importantly, the supervised step is not used to update the segmentation network, but to update an augmentation function ($A_{\boldsymbol{\theta}}$). This is made possible by backpropagating the supervised loss (2) through the update step (1).
1 Introduction
In medical imaging, the availability of high-quality reference data (high-resolution and artifact-free images, expert annotations, fair representation of populations) is a major limiting factor, compared to other computer vision domains [79]. This can be explained by the cost of data acquisition, noise and artifacts inherent to medical images, as well as the time and expertise required to label dense three-dimensional images. This limitation has hampered the development of modality-agnostic tools and led to a web of highly specific models trained on small datasets, whose performance degrades greatly in the presence of slight domain shifts [42, 78, 55], making out-of-site generalizability one of the most difficult and pervasive problems in biomedical machine learning [8]. This behavior, specific to supervised training on small datasets – peak on-domain performance with a rapid decay as data shifts away from the training domain – has long hindered the adoption of learning-based methods by the academic and clinical community. In contrast, supervised shallow models and unsupervised Bayesian approaches reach moderate but consistent performances across the board [66, 3].
In a data-limited regime, randomized augmentation remains crucial and provides performance boosts in most contexts [74, 67, 81]. The use of proper augmentation – along with methodical architecture and hyperparameter tuning – has been shown to be one of the main predictors of accuracy on an array of medical image segmentation challenges [2, 39]. The introduction of domain randomization [75, 5, 89, 76], in which the concept of augmentation is pushed even further by randomly sampling entire images, has been a paradigm shift in this regard. In its version most commonly used in biomedical imaging, domain randomization starts with a set of densely annotated shapes from which intensities are conditionally generated, followed by a set of common-but-extreme image augmentation steps [6]. This strategy ensures that labels and images are perfectly aligned by construction. Nonetheless, performance and generalizability are often sensitive to the choice of hyper-parameters that govern the statistical distributions of augmentations, which must be carefully validated on a dataset of real data. Furthermore, the synthesis procedure relies on ad hoc rules that may not capture the full range of features observed in real images, thereby setting a ceiling in achievable accuracy. Indeed, models trained with synthetic data always perform slightly worse than supervised models applied on-domain (but generalize significantly better) [6] due to the so-called “reality gap” [41].
Optimally, one would like to keep the best of both approaches, namely high on-domain accuracy and high generalizability to off-domain data. A simple solution consists of training models on a mix of synthetic and real data, sampled with some predefined proportion [59, 57]. However, it is possible to train a single network in such a way that it ends up containing multiple non-overlapping subnetworks [88], and it is likely that naively-trained deep models also internalize parallel subnetworks. Models trained on both synthetic and real data therefore risk overfitting some of their filters to the small set of real data. An alternative approach consists of generating realistic data from labels using losses that aim to match the distributions of synthetic and real images (adversarial [68, 28], contrastive [19, 31], normalizing flows [46, 61], diffusion [33, 58]). This however introduces an objective that is not relevant to the original task, leading to two major drawbacks. First, conditional labels and generated features are not constrained to be aligned, especially if the labels’ distribution deviates from the distribution of their target features in the real images [51]. Second, synthetic pairs may perceptually match real images without being able to train an accurate segmentation network. Indeed, distribution-matching models may spend too much capacity synthesizing features that have no impact on segmentation accuracy, or conversely fail to generate off-domain features that induce segmentation robustness.
Instead, we introduce Learn2Synth, a model trained with a single objective: improving the real world accuracy of a network trained on synthetic data (Figure 1). To do so, we proceed in two steps: (1) we freeze the synthesis network, push synthetic data through the segmentation network and update it; (2) we freeze the segmentation network, push real data through the segmentation network, and update the synthesis network. Naively, if a network is fed real data, no gradients flow back to the synthesis network; the key to our approach is to leverage hyper-gradients [53, 4, 25, 11], obtained by differentiating through step (1). In summary:
  1. 1.
    We propose a novel training pipeline for segmentation tasks where synthetic data, generated under ad hoc rules, is augmented by a trainable network. This avoids direct reliance on real data for training the segmentation network while still improving segmentation accuracy.
  2. 2.
    Instead of promoting realism using adversarial or perceptual loss functions, our strategy encourages the augmentation network to generate synthetic samples that optimize segmentation performance on real data.
  3. 3.
    By alternating between synthetic and real passes, the method improves the robustness and generalizability of the segmentation network, enhancing its performance across different image contrasts while preserving data-label pair integrity.
  4. 4.
    We demonstrate that our strategy improves performance across the board compared to pure domain randomization, while greatly improving generalizability compared to supervised training. Furthermore, the learned augmentation parameters give us insight into the optimal environment for network training.
2 Related work
Learning-based augmentation. Deep learning methods are prone to overfitting, especially when training data are limited. Data augmentation is a widely used technique to counteract this issue by artificially expanding the diversity of the training dataset. Traditional data augmentation techniques involve transformations such as intensity adjustments, spatial distortions, and geometric alterations [84]. These transformations help neural networks generalize better by exposing them to a wider range of variations. However, they rely on handcrafted transformations and hyperparameters that may not be optimal for specific datasets or tasks. Learning-based augmentation methods aim to alleviate this problem; they leverage deep learning models to learn optimal data transformations based on the dataset at hand, rather than relying on predefined parameters. One of the pioneering works in this area is AutoAugment [20], which uses reinforcement learning to identify the best augmentation policies for a given dataset, while Fast AutoAugment [50] and Faster AutoAugment [29] improve efficiency and scalability in the search process. In biomedical imaging, one typically has access to a large set of unlabeled data but a very small (or partial) set of labels. In this context, adversarial augmentation has been used to optimize augmentation parameters dynamically during training [85, 9, 14], producing challenging examples that encourage the model to learn more robust features. In fully unsupervised settings, contrastive learning has shown promise by generating multiple augmented views of the same image to capture invariant representations [10, 83], thereby leveraging unlabeled data to improve feature extraction and generalization. Alternatively, CycleGAN [90] and other simpler GAN-based methods have been used to translate labels to images and images to labels without requiring explicit pairing [70, 1].
Domain adaptation. A model trained on one domain usually does not perform well on images from another domain. Domain adaptation is a way to address the challenge and explicitly seeks to bridge the gaps between a source domain with labeled data, and a target domain without labels. One of the common solutions is to map both source and target domains to a common latent space [43, 22, 27, 82, 80]. By using image-to-image translation methods, generative adaptation methods have tried to match the source domains to the target domain [65, 37, 87]. Researchers have also proposed to operate in both feature and image space, leading to impressive results in cross-modality segmentation [13, 34]. For intra-modality scenarios, impressive performances have been obtained with test-time adaptation methods [44, 32], requiring only light fine-tuning at test-time.
Synthetic training data and domain randomization. Large amounts of high-quality labeled data are essential for the success of deep learning methods. It is always labor-intensive and costly to obtain high-quality annotations, especially for biomedical domains, which usually require domain knowledge and large amounts of work. Instead of using real data to train the networks, synthetic data can be introduced to increase generalization and robustness, either generated with physics-based models [42], adversarial generative networks [26, 12, 86, 49, 72, 77], or conditioned on label maps [54, 40, 6]. These strategies can generate large training datasets with perfect ground truth without any cost [63]. However, the generated synthetic images suffer from a “reality gap” [41]. In neuroimaging, domain-randomization underpins a family of contrast-invariant methods that target an array of tasks (SynthSeg [6, 7], SynthSR [38], SynthMorph [35], SynthStrip [36]). By randomizing the training data, networks can experience a virtually infinite variety of intensities and artifacts, which helps to reduce overfitting and enhance generalization to unseen data. However, this augmentation technique relies on ad hoc rules, which can constrain the diversity of shapes and artifacts generated.
3 Method
3.1 Training procedure
We have designed a training procedure in which synthetic training examples generated using ad hoc rules are enhanced by a trainable network, before being used to train a segmentation network. The segmentation network is trained only on synthetic examples, whereas the enhancement network is trained by backpropagating a loss computed only on real images (Figure 1). This strategy encourages the augmentation network to produce samples that improve the accuracy of the segmentation network on real data; leveraging real labeled examples without directly using them in training the segmentation network. In the following sections, $(\mathbf{x}_{\mathrm{real}},\mathbf{y}_{\mathrm{real}})$ and $(\mathbf{x}_{\mathrm{synth}},\mathbf{y}_{\mathrm{synth}})$ represent real and synthetic labeled examples, respectively, while $A_{\boldsymbol{\theta}}$ and $S_{\boldsymbol{\phi}}$ denote the augmentation and segmentation functions with corresponding weights $\boldsymbol{\theta}$ and $\boldsymbol{\phi}$. Our method consists of a synthetic pass and a real pass, which we describe in detail.
3.1.1 Synthetic pass
In the synthetic pass, a synthetic image, generated from a label map, is first pushed through an augmentation function $A_{\boldsymbol{\theta}}$. Its output is denoted as $A_{\boldsymbol{\theta}}(\mathbf{x}_{\mathrm{synth}})$, and the pair $(A_{\boldsymbol{\theta}}(\mathbf{x}_{\mathrm{synth}}),\mathbf{y}_{\mathrm{synth}})$ is subsequently used to train the segmentation network. The forward pass applied to the synthetic data yields
$\mathcal{L}_{\mathrm{synth}}\triangleq\operatorname{SoftDice}\left(S_{\boldsymbol{\phi}}\left(A_{\boldsymbol{\theta}}\left(\mathbf{x}_{\mathrm{synth}}\right)\right),\mathbf{y}_{\mathrm{synth}}\right),$ (1)
where $\operatorname{SoftDice}$ is the soft Dice loss function [71] and $\mathcal{L}_{\mathrm{synth}}$ represents the loss value. The gradient of this loss with respect to the segmentation network’s weights is computed by automatic differentiation and used to update the segmentation network through the optimizer’s update rule
$\mathbf{g}_{\phi}$ $\triangleq\frac{\partial\mathcal{L}_{\mathrm{synth}}}{\partial\boldsymbol{\phi}},$ $\boldsymbol{\phi}^{\ast}$ $\triangleq\operatorname{update}\left(\boldsymbol{\phi},\mathbf{g}_{\phi}\right).$ (2)
It is important to note that during the synthetic pass we only update the weights of the segmentation network ($S_{\boldsymbol{\phi}}$), while the parameters of the augmentation network ($A_{\boldsymbol{\theta}}$) remain fixed.
3.1.2 Real pass
Using only the synthetic pass, it is unsurprising that the trained segmentation network (with parameters $S_{\boldsymbol{\phi}}$) performs well on images generated by $A_{\boldsymbol{\theta}}(\mathbf{x}_{\mathrm{synth}})$. The next question is: how can we bridge the gap between $A_{\boldsymbol{\theta}}(\mathbf{x}_{\mathrm{synth}})$ and $\mathbf{x}_{\mathrm{real}}$ to ensure the segmentation network also performs well on $\mathbf{x}_{\mathrm{real}}$? Our solution is to leverage real-labeled data by introducing a real pass that encourages a change in $A_{\boldsymbol{\theta}}$ that increases the segmentation accuracy on $\mathbf{x}_{\mathrm{real}}$.
A real image $\mathbf{x}_{\mathrm{real}}$ is pushed through the (updated) segmentation network, resulting in the loss
$\mathcal{L}_{\mathrm{real}}\triangleq\operatorname{SoftDice}\left(S_{\boldsymbol{\phi}^{\ast}}\left(\mathbf{x}_{\mathrm{real}}\right),\mathbf{y}_{\mathrm{real}}\right).$ (3)
If implemented naively, no gradients would flow back into the augmentation network during the real pass because the real image does not pass through the augmentation network. To make this strategy effective, we implement a differentiable update step that encourages the augmentation network to generate synthetic images that improve the update of the segmentation network. The gradient of this loss with respect to the augmentation network is:
$\mathbf{g}_{\theta}\triangleq\frac{\partial\mathcal{L}_{\mathrm{real}}}{\partial\boldsymbol{\theta}}$ ${}=\frac{\partial\mathcal{L}_{\mathrm{real}}}{\partial\boldsymbol{\phi}^{\ast}}\times\frac{\partial\boldsymbol{\phi}^{\ast}}{\partial\mathbf{g}_{\phi}}\times\frac{\partial\mathbf{g}_{\phi}}{\partial\boldsymbol{\theta}}$
${}=\frac{\partial\mathcal{L}_{\mathrm{real}}}{\partial\boldsymbol{\phi}^{\ast}}\times\frac{\partial\boldsymbol{\phi}^{\ast}}{\partial\mathbf{g}_{\phi}}\times\frac{\partial^{2}\mathcal{L}_{\mathrm{synth}}}{\partial\boldsymbol{\phi}\partial\boldsymbol{\theta}^{T}},$ (4)
which is then used to update the augmentation network through the update step
$\boldsymbol{\theta}^{\ast}\triangleq\operatorname{update}(\boldsymbol{\theta},\mathbf{g}_{\theta})~{}.$ (5)
There are three components in the computation of $\mathbf{g}_{\theta}$:
  1. 1.
    $\frac{\partial\mathcal{L}_{\mathrm{real}}}{\partial\boldsymbol{\phi}^{\ast}}$ is the gradient of the real loss with respect to the current weights of the segmentation network. In a traditional training scenario using real data, this would be used to update the segmentation network’s weights.
  2. 2.
    $\frac{\partial\boldsymbol{\phi}^{\ast}}{\partial\mathbf{g}_{\phi}}$ is the gradient of the update step with respect to the previous gradient. If the update step is a gradient descent step, that term is simply an identity matrix scaled by the learning rate.
  3. 3.
    $\frac{\partial^{2}\mathcal{L}_{\mathrm{synth}}}{\partial\boldsymbol{\phi}\partial\boldsymbol{\theta}^{T}}$ is the Hessian of the synthetic loss with respect to both the segmentation and augmentation networks. This term indicates how the synthetic loss changes when taking steps in both $\boldsymbol{\phi}$ and $\boldsymbol{\theta}$.
Note that the Hessian is only introduced for interpretation purposes. In practice, the complete gradient is obtained by automatic differentiation; the Hessian matrix is therefore never allocated in memory.
3.2 Augmentation models
The “reality gap” that leads to a decrease in performance when applying a synthetically trained model to real data can have three different causes: (1) a miscalibration of the synthesis hyper-parameters, (2) overly simplistic synthesis rules that fail to generate realistic features, or (3) overly complicated synthesis results that force the network to waste capacity on features that never occur in real data. We introduce two learnable augmentation functions that target these possibilities:
  • a parametric procedure in which the hyper-parameters that govern Gaussian noise and intensity non-uniformity are learnable (Section 3.3);
  • a nonparametric procedure in which a learnable enhancement network is applied to synthetic data (Section 3.4).
3.3 Parametric model
Our parametric model encompasses two artifacts commonly found in magnetic resonance images: Gaussian noise and intensity non-uniformity (INU). For simplicity, we describe the procedure with one-dimensional signals. Its extension to two or three spatial dimensions is straightforward.
INU. Magnetic resonance images typically suffer from smooth changes in magnitude caused by the spatial profile of transmit and receive coils. While the transmit effect is nonlinear and sequence-dependent, the receive effect is multiplicative and commonly modeled using a small number of smooth basis functions. In order to make the spatial frequency of the receive field differentiable, we sample different random fields with different node spacings, which we modulate with learnable coefficients and combine. Formally, the generation of one of the random fields is written as
$\boldsymbol{\beta}_{k}$ $\sim U_{M_{k}}(0.5,2),$ $\boldsymbol{\alpha}_{k}$ $=\mathbf{B}_{k}\boldsymbol{\beta}_{k}$ (6)
where $M_{k}$ is the number of low-dimensional coefficients, $\mathbf{B}_{k}\in\mathbb{R}^{N\times M_{k}}$ is a first-order B-spline basis functions with node spacing $(N-1)/(M_{k}-1)$ and $N$ is the number of voxels in the image. The receive field and intensity-modulated images are then constructed by
$\boldsymbol{\alpha}$ $=\prod_{k=1}^{K}\boldsymbol{\alpha}_{k}^{c_{k}}~{},$ $\mathbf{x}_{\mathrm{synth}}$ $\leftarrow\mathbf{x}_{\mathrm{synth}}\odot\boldsymbol{\alpha}~{},$ (7)
where $\mathbf{c}\in\mathbb{R}^{K}$ are learnable coefficients, and products are taken element-wise.
Gaussian noise. Assuming that Gaussian noise variance is identical across all images in the dataset, we implement learnable noise augmentation by sampling Gaussian noise identically distributed across voxels and scaling it by a learnable standard deviation $\sigma$, resulting in the equation
$\boldsymbol{\varepsilon}$ $\sim\mathcal{N}_{N}(0,1),$ $\mathbf{x}_{\mathrm{synth}}$ $\leftarrow\mathbf{x}_{\mathrm{synth}}+\sigma\cdot\boldsymbol{\varepsilon}.$ (8)
However, the assumption of identical noise variance across images is very unlikely to stand in reality. Alternatively, we can learn a noise hyper-variance that is randomly modulated per-image:
$s\sim\mathcal{N}(0,1),~{}~{}~{}\boldsymbol{\varepsilon}\sim\mathcal{N}_{N}(0,1),$
$\mathbf{x}_{\mathrm{synth}}\leftarrow\mathbf{x}_{\mathrm{synth}}+\sigma\cdot s\cdot\boldsymbol{\varepsilon}.$ (9)
This step ensures that the model not only learns to generate realistic noise but also maintains the ability to handle various noise levels during the augmentation process.
Joint model. The set of learnable augmentation parameters is $\boldsymbol{\theta}~{}=~{}[\mathbf{c},\sigma]$. As described in Section 3.1, we use the $({\mathbf{x}_{\mathrm{synth}},\mathbf{y}_{\mathrm{synth}}})$ pairs to train the segmentation network $S_{\boldsymbol{\phi}}$, while the augmentation network $A_{\boldsymbol{\theta}}$ is responsible for generating diverse synthetic images.
3.4 Nonparametric model
Rather than relying on a learnable parametric model in the synthesis pass (such as setting the amplitude of uniform noise), we route the standard synthetic output through a UNet. The output of this UNet is then added to the original input, effectively learning a residual that optimizes the synthetic data. This approach eliminates the need to manually define parametric models for each type of augmentation. Such a model would however be deterministic, whereas we aim to implement a randomization network. Inspired by diffusion networks, we concatenate the input image with a channel of Gaussian noise, thereby providing additional entropy to the network, leading to the following formulation:
$\boldsymbol{\xi}$ $\sim\mathcal{N}_{N}(0,1),$ (10)
$\mathbf{x}_{\mathrm{synth}}$ $\leftarrow\mathbf{x}_{\mathrm{synth}}+A_{\boldsymbol{\theta}}([\mathbf{x}_{\mathrm{synth}},\boldsymbol{\xi}])~{}.$ (11)
Furthermore, in order to obviate forcing the augmentation network to learn an identity mapping, we apply the parametric noise augmentation described in Equation 9 to its output.
4 Experiments
Our experiments aim to answer two fundamental questions: (1) what training environment is deemed optimal by Learn2Synth when the domain gap between $\mathbf{x}_{\mathrm{real}}$ and $\mathbf{x}_{\mathrm{synth}}$ is known? (2) can Learn2Synth bridge the on-domain performance gap with supervised approaches without paying a penalty off-domain? We tackle (1) in Section 4.1 using synthetic experiments, in which $\mathbf{x}_{\mathrm{real}}$ is defined as a stochastic function $A_{{\hat{\boldsymbol{\theta}\mkern 2.0mu}\mkern-2.0mu}{}}(\mathbf{x}_{\mathrm{synth}})$ with known parameters ${{\hat{\boldsymbol{\theta}\mkern 2.0mu}\mkern-2.0mu}{}}$. The parameters of that same function $A_{\boldsymbol{\theta}}$ are then optimized using our framework, and the optimal parameters ${\boldsymbol{\theta}}^{\ast}$ compared to the true parameters ${{\hat{\boldsymbol{\theta}\mkern 2.0mu}\mkern-2.0mu}{}}$. We tackle (2) in Section 4.2 using two experiments with real MRI data, that mimic two typical training strategies for brain segmentation networks. In the first experiment, we use a single imaging contrast and a large set of “soft truth” labels obtained with an automated method. In the second experiment, we use a small number of high-quality “ground truth” labels and investigate generalizability across contrasts.
Implementation details. All experiments use 2D coronal slices extracted from 3D brain MR images. We use the cornucopia package [^1] https://github.com/balbasty/cornucopia as our synthetic generator and use a standard UNet [64] as the backbone for our segmentation and nonparametric augmentation networks. Architecture details are provided as supplementary material. The networks are randomly initialized and trained from scratch. We use the soft Dice loss [71] to supervise the training of the segmentation network. We use the Adam optimizer [45] with a learning rate of $1\times 10^{-3}$.
4.1 Experiments on synthetic datasets
Dataset. We automatically segmented 434 images from the OASIS dataset [48], using FreeSurfer [23] and used the resulting label maps throughout. We used 60% of the samples as the training set, 20% as the validation set, and the remaining 20% as the test set. At each iteration, noise-free synthetic images are generated from these label maps by assigning random intensities to each label. New images are generated at every epoch, yielding a virtually infinite number of synthetic pairs. Overall results are therefore minimally affected by the portion of training samples.
Settings. Here, synthetic images ($\mathbf{x}_{\mathrm{synth}}$) are noise-free, while “real” images ($\mathbf{x}_{\mathrm{real}}$) are obtained by adding noise and/or INU to noise-free images, with predefined parameter ${{\hat{\boldsymbol{\theta}\mkern 2.0mu}\mkern-2.0mu}{}}$. A Learn2Synth model is then trained on this joint dataset, using either the parametric or nonparametric augmentation model $A_{\boldsymbol{\theta}}$. In each sub-experiment, we vary the value of the true parameter ${{\hat{\boldsymbol{\theta}\mkern 2.0mu}\mkern-2.0mu}{}}$ and compare it with the value of the learned parameter ${\boldsymbol{\theta}^{\ast}}$. In the parametric case, if ${\boldsymbol{\theta}^{\ast}}$ converged to ${{\hat{\boldsymbol{\theta}\mkern 2.0mu}\mkern-2.0mu}{}}$, we would effectively obtain the same segmentation network as if $S_{\boldsymbol{\phi}}$ had been directly trained on $\mathbf{x}_{\mathrm{real}}$. However, there is no built-in incentive for Learn2Synth to recover the exact same parameters. Instead, it is solely tasked with optimizing segmentation performance on $\mathbf{x}_{\mathrm{real}}$.
Baselines. In each sub-experiment, we train two SynthSeg baselines: one where the training data uses the same augmentation parameters as the “real” data (Naive SynthSeg), and one where it uses the parameters inferred by Learn2Synth (Optimized SynthSeg). In cases where the inferred parameters ${\boldsymbol{\theta}^{\ast}}$ differ from the “real” parameters ${{\hat{\boldsymbol{\theta}\mkern 2.0mu}\mkern-2.0mu}{}}$, these baselines allow us to verify that the inferred parameters are indeed optimal.
4.1.1 Gaussian noise
We use the homogeneous noise model described in Equation 8, with different values of $\sigma=\{0,0.05,0.1,0.15\}$, as well as a version where $\sigma$ is randomly sampled in $U[0.025,0.2]$ at each step. In all cases, the learnable augmentation model $A_{\boldsymbol{\theta}}$ implements Equation 8. We report the inferred $\sigma$s for different noise-level settings in Table 1. Inferred parameters match well with the preset parameters, although we note that the optimal solution could have been a different set of parameters in which the optimal training noise levels might be larger (or smaller) than those in the test images, should they yield higher real-world accuracy.
Preset $\hat{\sigma}$ 0 0.050 0.100 0.150 [0.025 0.2]
Inferred $\sigma^{\ast}$ 0.001 0.042 0.098 0.146 0.134
Table 1: Inferred noise $\sigma$ under the “noise-only” setting.
In Table 2, each row denotes the segmentation accuracy of models trained on different levels of noise when tested on images with some other level of noise. For example, the first block ($\sigma=0$) shows the accuracy of models trained on noise-free data when tested on images with five different noise levels. Conversely, each column contains the results of all models, when tested on images with a given amount of noise. We highlight the highest (bold) and second highest (underline) numbers in each column. As expected, performance declines as the noise level increases in the test set. Additionally, for images subjected to a specific noise level, the model trained under the same noise conditions achieves optimal performance. This is evidenced by the highlighted diagonal. This finding is in line with results from Table 1. Most importantly, when tested on noisy images under models trained with matching noise conditions, Learn2Synth consistently outperforms naive SynthSeg. In theory, after convergence, Learn2Synth and Optimized SynthSeg should yield identical results, which is also verified in practice.
Train Test Model $\sigma$ = 0 $\sigma$ = 0.05 $\sigma$ = 0.1 $\sigma$ = 0.15 $\sigma\sim U$
$\sigma$ = 0 N 0.913 0.310 0.209 0.142 0.205
L 0.923 0.346 0.172 0.092 0.151
O 0.919 0.361 0.189 0.105 0.167
$\sigma$ = 0.05 N 0.889 0.874 0.718 0.520 0.646
L 0.900 0.887 0.744 0.492 0.635
O 0.898 0.890 0.736 0.508 0.651
$\sigma$ = 0.1 N 0.867 0.858 0.841 0.772 0.815
L 0.874 0.873 0.858 0.799 0.831
O 0.871 0.869 0.847 0.810 0.829
$\sigma$ = 0.15 N 0.864 0.857 0.847 0.829 0.838
L 0.860 0.855 0.854 0.840 0.834
O 0.867 0.860 0.852 0.836 0.841
$\sigma\sim U$ N 0.870 0.867 0.850 0.826 0.831
L 0.866 0.864 0.851 0.833 0.839
O 0.872 0.870 0.853 0.837 0.843
Table 2: Segmentation accuracy (Dice) for naive SynthSeg (N), Learn2Synth (L) and Optimized SynthSeg (O).
4.1.2 Intensity non-uniformity
Here, we use both the INU model described in Equation 7 and the homogeneous noise model described in Equation 8. For the former, we use $K=3$ random fields with increasing spatial frequency ($\left\{M_{1},M_{2},M_{3}\right\}=\left\{2,4,8\right\}$). We denote the corresponding learnable parameters $c_{\mathrm{low}}$, $c_{\mathrm{mid}}$ and $c_{\mathrm{high}}$, with preset values $\hat{c}_{\mathrm{low}}=\hat{c}_{\mathrm{mid}}=\hat{c}_{\mathrm{high}}=0.5$. Noise-free target “real”, and learned images are shown in Figure 2. Inference results for the value of the noise parameter are reported in Table 3, and are in line with those obtained in the noise-only experiment.
Refer to caption Refer to caption Refer to caption
(a) Noise-free image (b) Learned image (c) Real image
Figure 2: Illustration of learned synthesized image. (a) Noise-free image, (b) learned synthesized image, and (c) real image.
Preset $\hat{\sigma}$ 0 0.050 0.100 0.150 [0.025 0.2]
Inferred $\sigma^{\ast}$ 0.018 0.049 0.085 0.147 0.134
Table 3: Inferred noise $\sigma$ under the “noise and INU” setting.
Similar to the “parametric noise” setting, in addition to examining the inferred noise levels, we also report the segmentation accuracy for various noise-level settings in Table 4 for the Learn2Synth model. We highlight the highest (bold) and second highest (underline) numbers in each column as well. Notably, both of the key observations from Section 4.1.1 hold true for the scenario where both the noise and the bias field are modeled. This continued consistency demonstrates that our approach performs well under varying conditions, whether noise alone or both noise and bias fields are considered.
Model Train Test $\sigma$ = 0 $\sigma$ = 0.05 $\sigma$ = 0.1 $\sigma$ = 0.15 $\sigma\sim U$
Parametric $\sigma$ = 0 0.924 0.867 0.748 0.619 0.729
$\sigma$ = 0.05 0.900 0.894 0.867 0.696 0.744
$\sigma$ = 0.1 0.894 0.889 0.872 0.832 0.849
$\sigma$ = 0.15 0.881 0.869 0.856 0.854 0.862
$\sigma\sim U$ 0.875 0.868 0.870 0.851 0.867
Nonparametric $\sigma$ = 0 0.903 0.871 0.752 0.603 0.769
$\sigma$ = 0.05 0.891 0.886 0.859 0.749 0.792
$\sigma$ = 0.1 0.884 0.877 0.865 0.828 0.839
$\sigma$ = 0.15 0.878 0.864 0.851 0.837 0.844
$\sigma\sim U$ 0.869 0.858 0.843 0.831 0.850
Table 4: Segmentation accuracy for Learn2Synth under the parametric “noise and INU” and nonparametric settings.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Input (b) Ground Truth (c) Supervised (d) SAMSEG (e) Naive SynthSeg (f) Learn2Synth
Figure 3: Qualitative results on OASIS3 and ABIDE. (a), (b) show the input and segmentation. (c)-(f) show the segmentations of different methods.
We have also trained Equation 9’s “nonparametric + noise” model, instead of a “INU + noise” model, on the same training data (i.e., generated using the parametric INU + noise function). While inference results are in line with previous sections (not shown here), segmentation accuracy is slightly lower than that obtained with an accurate parametric model. This highlights that if a low-dimensional parametric model of the data is known, it should be used over a high-dimensional nonparametric alternative.
4.2 Experiments on real-world datasets
In the previous section, we have shown that using synthetic data Learn2Synth is able to transform its synthetic training domain so that it matches its “real” training domain. Furthermore, when domains are not matched exactly, the resulting network is more accurate than a network trained on the “real” domain. However, the “real” domain that we used was far from realistic, and the reality gap between typical synthetic data and real MR images is much wider. In this section, we aim to demonstrate the performance of Learn2Synth in real world scenarios.
Our first experiment (Section 4.2.1) aims to show that even a well-specified synthesis model can be improved if real data is available. We use images from two large multi-site datasets, along with “soft truth” labels obtained with a standard automated method.
Our second experiment (Section 4.2.2) tackles a more interesting problem, where a large set of “soft truth” labels is available, but only a handful of manual “ground truth” labels are. In this context, we further investigate the performance of our method and its baselines on a held-out test set whose contrast differs from the one used during training.
4.2.1 Learn2Synth outperforms mixed training
Datasets. We automatically segmented anatomical images from the ABIDE [21] and OASIS3 [48] datasets, using FreeSurfer [23] and merged the resulting segmentation maps into four labels (white matter, cortical gray matter, subcortical gray matter, and ventricles). We used 60% of the samples as the training set, 20% as the validation set, and the remaining 20% as the test set. Both datasets are multicentric; ABIDE aimed to study participants in the autism spectrum, with an age range slightly biased towards younger participants; OASIS aimed to study patients with dementia, with an age range slightly biased towards older participants. Both datasets mostly contain images acquired with a MPRAGE sequence, which is T1-weighted and optimized for cortical contrast.
Baselines. We compare Learn2Synth with a simple Supervised UNet [64] trained on real data; SAMSEG [62], a state-of-the-art unsupervised Bayesian segmentation framework; and SynthSeg [6], a state-of-the-art domain randomization approach for brain image segmentation. Note that since we are running our experiments in 2D, we do not use the published SynthSeg model, but instead train, from scratch, our own version with matched hyper-parameters. We call this baseline Naive SynthSeg for clarity. We also train another SynthSeg model on a dataset augmented with a small number of real images (Mixed SynthSeg), and we further finetune Naive SynthSeg on the same small number of real images (Finetuned SynthSeg).
Method ABIDE OASIS3
Supervised UNet 0.908 0.899
SAMSEG 0.875 0.841
Naive SynthSeg 0.869 0.831
Mixed SynthSeg 0.875 0.854
Finetuned SynthSeg 0.871 0.847
Learn2Synth (parametric setting fixed $\sigma$) 0.878 0.860
Learn2Synth (parametric setting varying $\sigma$) 0.881 0.857
Learn2Synth (nonparametric setting fixed $\sigma$) 0.879 0.875
Learn2Synth (nonparametric setting varying $\sigma$) 0.874 0.881
Table 5: Segmentation accuracy for Learn2Synth and the baselines on ABIDE and OASIS3.
Figure 4: Results obtained when training on MPRAGE scans and testing on FLASH scans.
Results. Qualitative segmentation results are presented in Figure 3 and quantitative results are summarized in Table 5Figure 3 shows that Learn2Synth produces cleaned and more accurate segmentation results compared to SAMSEG and naive SynthSeg, which is further validated by the higher Dice scores reported in Table 5. Learn2Synth also obtains higher Dice scores than the two SynthSeg variants that incorporate real data. These results highlight the effectiveness of Learn2Synth in enhancing segmentation quality for complex real-world data.
Setting # of Train, Val., Test Test Set (MPRAGE) $3^{\circ}$ $5^{\circ}$ $20^{\circ}$ $30^{\circ}$
SynthSeg / $0.861\pm 0.028$ $0.776\pm 0.012$ $0.694\pm 0.027$ $0.766\pm 0.018$ $0.781\pm 0.016$
Supervised UNet 29, 5, 5 $\mathbf{0.941}\pm 0.002$ $0.419\pm 0.025$ $0.396\pm 0.020$ $0.671\pm 0.026$ $0.769\pm 0.022$
Supervised UNet 5, 5, 29 $0.907\pm 0.007$ $0.397\pm 0.018$ $0.413\pm 0.016$ $0.586\pm 0.033$ $0.692\pm 0.033$
Supervised UNet 1, 5, 33 $0.885\pm 0.010$ $0.267\pm 0.019$ $0.247\pm 0.013$ $0.544\pm 0.026$ $0.651\pm 0.025$
Learn2Synth 29, 5, 5 $0.895\pm 0.011$ $\mathbf{0.804}\pm 0.014$ $\mathbf{0.789}\pm 0.010$ $0.785\pm 0.025$ $0.797\pm 0.023$
Learn2Synth 5, 5, 29 $0.867\pm 0.030$ $0.798\pm 0.013$ $\mathbf{0.789}\pm 0.010$ $\mathbf{0.795}\pm 0.026$ $\mathbf{0.799}\pm 0.024$
Learn2Synth 1, 5, 33 $0.718\pm 0.045$ $0.606\pm 0.020$ $0.603\pm 0.017$ $0.690\pm 0.036$ $0.707\pm 0.032$
Table 6: Performance comparison of different models across various datasets.
4.2.2 Learn2Synth generalizes to unseen contrasts
Datasets. We further validate the proposed algorithm on Buckner39 [23], a dataset with 39 manually labeled MPRAGE images, and the Freesurfer maintenance dataset [24], which contains images acquired in 8 subjects with a FLASH sequence at multiple flip angles. In contrast with MPRAGE, the FLASH sequence does not apply an inversion pulse, which results in lower cortical contrast. Furthermore, different flip angles give rise to different contrasts. We use FLASH scans acquired with flip angles of $3^{\circ}$ and $5^{\circ}$ (proton density-weighted) and $20^{\circ}$ and $30^{\circ}$ (T1-weighted), with the latter being closer in appearance to MPRAGE scans than the former. All scans were skull-stripped and manually delineated with the same protocol as in [23]. The synthetic portion of the training set used label maps derived from the ABIDE and OASIS3 datasets.
Settings. We trained a SynthSeg baseline using the same settings as in the previous section. We also trained three supervised UNets and three Learn2Synth models on the Buckner39 dataset, with different training set sizes. Varying the training size allows different scenarios of data availability to be modeled. All models were tested on the remaining test portion of Buckner39 and on all images from the Freesurfer maintenance dataset. Models that obtained the best performance on their validation set were selected for evaluation on the test set. To quantify the impact of this validation step, we also provide results obtained with fully converged models in supplementary material.
Results. Quantitative results are summarized in Table 6, and graphically represented in Figure 4. The supervised UNet achieves the highest Dice score on Buckner39, particularly when trained with 29 training samples, reaching $0.941\pm 0.002$, which is not surprising. However, its performance drops significantly on FLASH scans, highlighting its limited generalization ability. Learn2Synth, on the other hand, demonstrates improved generalization across datasets, achieving Dice scores of approximately 0.79 across all FLASH sequences when trained with as few as 5 training samples. Surprisingly, Learn2Synth improves its Dice over the SynthSeg baseline on MPRAGE scans (as expected) but also on all FLASH scans. These results suggest that Learn2Synth effectively generalizes to unseen imaging conditions.
5 Conclusion
In conclusion, we present Learn2Synth, a novel approach to enhance segmentation performance without directly exposing the segmentation network to real data. Our experimental results on both synthetic and real datasets confirm that Learn2Synth reaches substantially higher segmentation performance accuracy than naive SynthSeg, and generalizes considerably better to unseen contrasts than supervised models as well as SynthSeg.
Acknowledgements. YB is supported by a fellowship from the Royal Society ($\|$$\|$232460). Support for this research was provided in part by the BRAIN Initiative Cell Census Network grants U01MH117023 and UM1MH130981, the Brain Initiative Brain Connects consortium (U01NS132181, 1UM1NS132358-01), the National Institute for Biomedical Imaging and Bioengineering (1R01EB023281, R21EB018907, R01EB019956, P41EB030006), the National Institute on Aging (R21AG082082, 1R01AG064027, R01AG016495, 1R01AG070988), the National Institute of Mental Health (UM1MH130981, R01 MH123195, R01 MH121885, 1RF1MH123195), the National Institute for Neurological Disorders and Stroke, (1U24NS135561-01, R01NS070963, 2R01NS083534, R01NS105820, R25NS125599), and was made possible by the resources provided by Shared Instrumentation Grants 1S10RR023401, 1S10RR019307, and 1S10RR023043. Additional support was provided by the NIH Blueprint for Neuroscience Research (5U01-MH093765), part of the multi-institutional Human Connectome Project. Much of the computation resources required for this research was performed on computational hardware generously provided by the Massachusetts Life Sciences Center (https://www.masslifesciences.com/) In addition, BF is an advisor to DeepHealth, a company whose medical pursuits focus on medical imaging and measurement technologies. BF’s interests were reviewed and are managed by Massachusetts General Hospital and Partners HealthCare in accordance with their conflict of interest policies.